The current price of a non-dividend-paying stock is $100. Over the next year the stock is expected either to rise to $110 or to fall to $90. An investor buys two put options with a strike price of $105.

Which of the following is necessary to delta-hedge the position?

A. Buy 0.5 shares of the stock.

B. Sell 0.5 shares of the stock.

C. Buy 0.25 shares of the stock.

D. Sell 0.25 shares of the stock.

E. None of the above.
Answer #1

Delta of Put=(MAX(105-110,0)-MAX(105-90,0))/(110-90)=-0.75

To delta hedge, 0.75*2=1.5 shares need to be bought

Option E